If the reserve ratio is 10 percent and a bank receives a new deposit of $10, this bank

a.

must increase required reserves by $1.

b.

will initially see its total reserves increase by $1.

c.

will be able to make new loans up to a maximum of $1.

d.

All of the above are true.
Answer #1

Answer

Option a

must increase required reserves by $1

==========

required reserves =deeposit * required reserve ratio

=10*0.1

=$1

=======

Excess reserves =new deposits -required reserves=10-1=9

the reserves increase by $10 and with required reserves of $1 and excess reserves of $9 where the bank can lend $9 the excess reserves.